Find the equation of the lines which pass through the point of intersection of the lines 4x-3y-1=0 and 2x-5y+3=0 and are equally inclined to the axes. 

Grade:11

1 Answers

sunil raikwar
9 years ago
Intersection point of 4x-3y-1=0 & 2x-5y+3=0 is (1,1).
if lines are equally inclined with the axes then slope of lines either 1 or -1.
so equations of lines are
y-1=1(x-1) & y-1=-1(x-1)
 
Thanks & Regards
Sunil Raikwar
askIITians Faculty